# Retrieving snow depth distribution by downscaling ERA5 Reanalysis with ICESat-2 laser altimetry

### v1

@@ 5.2 Snow depth downscaling
Tree-structure models perform well in capturing nonlinear relationships, and the model used in this study provides reasonable predictions for snow depth. However, unless calibration is applied, our approach leaves a significant amount of variability unaccounted for, particularly for periods/areas with high snow depths as observed in the time series analysis (Figure 7) and in the snow depth maps (Figures S.8, S.9). The underestimation of spatial variability arises from several factors: i) the regression model’s inherent limitations, ii) a likely imbalance of the training samples, and iii) the limitation of the predictors to capture physical processes driving the redistribution of snow.
@@
Inherently (i), deterministic regression models primarily yield conditional means and may not adequately represent extreme events, especially when key features do not fully explain variations. This often results in underfitting as the models strive to minimize overall prediction errors. Additionally (ii), the spatial distribution of snow depth is inherently scale-dependent [55, 60], meaning a sensor’s spatial resolution is critical in determining captured variability. Our training samples, derived from the ICESat-2 ATL08 elevation product, correspond to an area of ca. 20 m × 15 m, which is better than previous study cases but still averages out finer-scale variability. ICESat-2’s sparse spatial sampling pattern with several kilometres gap between sampled snow depth profiles, combined with a 92-day revisit period of the satellite and possible cloud cover leads to substantial gaps in seasonal data coverage. Data points with high snow depth measurements are thus likely to be relatively few. ICESat-2’s elevation data, derived from a ground-finding algorithm that uses a probability distribution function (PDF) of reflected photons [67], can vary in quality depending on the terrain. In rugged terrain where snow depth is typically the deepest, noise or unbalanced sampling can lead to misrepresentation of extreme values.
@@
The shortcomings of the model and data sample could be mitigated if model predictors were able to fully capture spatial variability of snow and underlying physical processes (iii), which is not the case. ERA5 Land input snow depth values may provide incorrect prior information as the dataset seems to contain interpolation effects that are possibly a downscaling artefact from the production of ERA5 Land from the coarser ERA5 data [32]. For example, snow depth values are much higher for ERA5 Land grid cells on or adjacent to large glaciers like the Jostedalsbreen ice cap in Southern Norway, and lower for grid cells on or adjacent to large water area. Consequently, although we have excluded all glaciers and lakes, neighbouring ERA5 Land grid cells still contain these biasing effects. There, the predicted subgrid variability is not solely a reflection of inherent snow conditions but is significantly influenced by this spatially-dependent bias that the model needs to detect and correct for based on spatial information,i.e. coordinate features (N, E). This underscores the importance of dense, localized sampling of ICESat-2 snow depths, which may be challenging for smaller study areas and for lower latitudes, given ICESat-2’s sparse sampling pattern. By expanding the training area, as done in this study, the model benefits from more samples but might also learn a more generalized, averaged representation of snow depth. Among the predictor features elevation emerges as the most significant, as expected from the temperature–elevation relationship, orographic and precipitation shadowing effects [60, 71] where also aspect plays an important role. The TPI at a 90-meter scale is a strong predictor (Figure 9) and corresponds to the size of terrain features in our study area. However, these predictors represent general patterns only and they are not sufficient to represent locally different snow depth. Here, the cumulative effect of multiple snow transport events over an entire winter season plays a crucial role. By refining the wind-aspect factor [2, 15] into new accumulative and erosion values to describe wind redistribution across topographic barriers, our model is able to explain a considerable amount of the observed preferential snow deposition (Figure 9 g). Erosion processes are not equally well captured (Figure 9 f), and the downscaling model still fails to correctly reproduce the net effect of the processes at play. There is potential to improve wind-related predictors, as the spatial-temporal resolution of wind fields from ERA5 Land monthly data is likely insufficient for capturing complex wind-topography interactions. Downscaling of wind fields [19] could enhance the model’s ability to capture wind-driven variability. Additionally, the model does not account for the impact of wind on the energy balance at the snow-atmosphere interface, involving sensible and latent heat exchanges [59], a factor that plays a crucial role in snow dynamics. The choice of predictors in this study is based on our expertise and corresponds to commonly used predictors in statistical modelling of snow distribution, but this choice inherently introduces bias as predictors are not able to explain all variability. Future research could benefit from incorporating more features that are related to physical snow processes or the integration of snow melt-out information from fractional snow cover data from optical satellite imagery [22] commonly used in snow data assimilation schemes with process-based models [51, 52]. Used as an exploratory tool, the model might assist in learning what processes are at play and how to best represent them statistically, ideally leading to a simplification of the workflow in the future.
@@
The introduced calibration step uses quantile mapping based on the observed cumulative snow depth distribution function during 2009 peak snow conditions in the Hardangervidda area. Our selected control points cover the entire study area (~1.5 points per square kilometre), ensuring a comprehensive correction across all regions analyzed. While greatly improving the results, especially for the reproduction of extremes values, the need of a scaling calibration in our approach is a disadvantage for a global application of our method. The slight underestimation of snow depths for the presented western sub-region in the validation area (Figure 6) suggests that the single scaling function obtained from the Hardangervidda area does not fit the sub-region perfectly, and a locally adjusted scaling might result in a better fit. Future research is needed to determine optimal training/model area sizes or transferable scaling calibration functions for areas with different topography and climate. However, extensive, other publicly available snow depth reference datasets like the Hardangervidda ALS dataset are currently not available for Norway and are generally very rare, as snow depth surveys are expensive and often connected to for-profit activities such as hydropower. In western North America, the Airborne Snow Observatory data [ASO; 70] includes multi-temporal snow depth maps of several large catchments that could be suitable to examine the local dependency of our downscaling approach and further develop the method. As an alternative to using extensive validation data for model result calibration, future research could explore the use of small representative snow depth distributions such as strategically chosen, smaller snow depth maps from drone-based surveys, data from snow courses, or snow depth measurements from meteorological stations.
@@ 5.3 Validation
Snow depth distribution in alpine environment is influenced by distinct processes occurring at various spatial and temporal scales. Our model demonstrates high fidelity of snow distribution patterns even at the microscale (Figure 6 c), though with lower performance metrics compared to the mesoscale. Our findings complement the insights from Mott et al. [60], which recommend a resolution finer than 50 meters to capture the spatial variability inherent in wind-driven processes. For example, Trujillo et al. [84], Mott et al. [58], Mendoza et al. [56] report a distinct ’scale break’ at around 100 m, with a stronger spatial autocorrelation of snow depths below the scale break than beyond. Such spatial correlation can be attributed to the wavelength of wind-driven snow accumulation features [60]. The ALS data and our model output identified the presence of scale breaks and dependence on terrain features (95 m accounting for 79% variance for ALS data and 108 m for 60% variance for our model, Figure 10), demonstrating the effectiveness of our scheme in capturing a high level of spatial details and the impact of landscape patterns. Notably, we also found a scale break at 1,300 m by fitting the empirical variograms at two different ranges, which corresponds to the wavelength of ridges in this area (Figure 10). These scale breaks underscore the critical resolution and minimal survey scope required to accurately capture spatial variability over the hillslope. Such insights are pivotal for designing snow surveys or evaluating snow models/products.
RMSE or $R^{2}$ are standard metrics for evaluating regression model performance, focusing primarily on overall accuracy and fit. However, they may not adequately capture the variability and structural patterns in snow depth distribution, especially at the microscale. Figure 6 and 7 show high signal fidelity (high Spearman rank correlation) however with poor $R^{2}$. To enhance the evaluation of regression outcomes for snow depth subgrid variability, it is essential to consider the scales involved and incorporate additional metrics. As an alternative to the Spearman rank correlation used in this study, structural similarity index measure (SSIM, Parr et al. 71) might be a suitable metric to capture relative similarities.
@@ 5.4 Application
Current limitations of ICESat-2 data, including the month-long data availability lag and significant spatial and temporal gaps, restrict its utility [13]. Spatio-temporal propagation of ICESat-2 snow depths by use of modelling is thus the logical next step to produce timeseries of snow depth maps and make this data useful for wider applications. Hereby, standard statistical models that solely depend on topographic features for snow spatial distribution fall short on temporal variability. Our workflow resolves many of these constraints by incorporating temporal variability from ERA5 Land. Other emerging approaches to create snow depth maps from ICESat-2 data include combining ICESat-2 snow depths with snow cover information and a process-based model within a data assimilation framework (Mazzolini et al. 52, using higher-resolution snow depths from ATL03 data) or a combination of deep learning and probabilistic data assimilation methods (Guidicelli et al. 26, using snow depth tracks synthesised from snow depth maps to mimic ICESat-2 snow depths). Both approaches yield promising results and also use ERA5 Land data as forcing, thus are in principle globally applicable, but focus on smaller catchments and coarser spatial resolution than presented in this work. Future research could focus on comparing or merging these different approaches.
Notably, we find that the performance of aggregated downscaled snow depth is not sensitive to the absolute precision of reference DEMs (Supplementary Section D), indicating that our approach is not limited to areas with high-resolution DEMs but can be applied using global DEMs in treeless conditions globally. However, careful consideration is required when applying this workflow to new study areas or applications: (1) global DEMs like the Copernicus GLO-30 or FABDEM are a patchwork of data from satellite-derived DEMs and local/national data sources and may not have equally good quality in different regions. (2) Care has to be taken to ensure including enough ICESat-2 samples in/around the target area. Areas further south will have larger coverage gaps due to ICESat-2’s sampling pattern that is densest at the poles. Upcoming ICESat-2 data expected for several more years will mitigate this. (3) applying calibration with the most representative field survey. (4) Accounting for regional climatic variability that some area are heavily influenced by specific weather events or patterns.
@@
The workflow described herein is applicable to a range of applications in need to assess snow distribution patterns where sparse snow observations exist. For example, our workflow can (1) mitigate data gaps in remote areas spatially and temporally, e.g. by interpolating and extrapolating weather station observations (Figure 6, 7), (2) help to plan and reduce cost of new snow surveys (e.g coordinating surveys with ICESat-2 coverage), (3) improve local estimates of snow water equivalent in large alpine watershed, and (4) be combined with a classical snow modeling approach. Suppose a snow model provides not just average value snow depth, but also a probability density distribution representing the subgrid varaibility [23], such a curve can then be translated into a snow depth map using our snow downscaling model.
@@
Given careful calibration and information on snow density, the presented cost-efficient subgrid parameterization for snow depth could be used to estimate snow water equivalent (SWE) and serve as input for discharge modeling [31]. Microscale snow depth maps can benefit studies of ecosystems in snow-covered regions, from habitat availability for wildlife to plant phenology under the snowpack. Snow depth parametrizations/distributions at the mesoscale for large areas or detailed snow depth maps at the microscale are crucial for local permafrost studies [23].
@@ 6 Conclusion
This study introduces a workflow for snow depth retrieval from ICESat-2 ATL08 and DEM data to downscale ERA5 Land snow depth data using an XGBoost tree-structure machine learning model. The two datasets, ICESat-2 and ERA5 Land, have complementary resolutions in space and time that allow for the generation of accurate monthly snow depth maps at the hillslope scale. Hereby, ERA5 Land data primarily provides the temporal variability, and ICESat-2 the spatial variability in snow depths, propagated in space using terrain features and other relevant predictors to train the downscaling model. Advanced bias correction and calibration are part of the workflow to address inherently present systematic errors in the data and correct for residual bias.
@@
    There are very few snow depth observations available in remote areas, and to date, no inexpensive ways to map small-scale variability exist. Here, ICESat-2 ATL08 data as presented in our workflow of co-registration and bias-correction stands out as a valuable data source.
@@
    The downscaling-calibration scheme’s performance to predict peak snow for 6×2 ALS flight strips in the Hardangervidda is very good at mesoscale (100 x 500 m, $R^{2}$ values ranging from 0.74 to 0.88). At microscale (10 m), the spatial snow depth pattern is captured very well but absolute values are less so. The model is also able to represent the spatial pattern of snow melt-out during late spring as visible from snow cover satellite data.
@@
    Post-calibration of downscaling model results is currently necessary to compensate for the model’s under-representation of extreme values. This under-representation is likely caused by inherent model behaviour (tending towards average values), remaining bias and the nature of the training data. Biasing factors include persistent snow patches on supposedly snow-free reference DEMs, simplified representation of wind processes, undersampling and underestimation of high snow depths in ICESat-2 data, and temporal nonstationarity in ERA5 Land data. Future research and more adaptable calibration methods for varied scenarios may improve results and remove the need of post-calibration.
@@
This work was supported by the SNOWDEPTH project, Research Council of Norway (Grant number 325519) and the research project ClimLAND, an EEA/EU collaboration grant between Norway and Romania 2014-2021, project code RO-NO-2019-0415,1290 contract no. 30/2020.

### v2

@@ 5.2 Snow depth downscaling
Tree-structure models perform well in capturing nonlinear relationships. The model used in this study provides reasonable predictions for snow depth. However, unless calibration is applied, our approach leaves a significant amount of variability unaccounted for, particularly for periods/areas with deep snow depths as observed in the time series analysis (Figure 7) or in the snow depth maps (Figures S.8, S.9). The underestimation of spatial variability arises from several factors: i) the regression model’s inherent limitations, ii) a likely imbalance of the training samples, and iii) the limitation of the predictors to capture physical processes driving the redistribution of snow.
@@
Inherently (i), deterministic regression models primarily yield conditional means and may not adequately represent extreme events, especially when key features do not fully explain variations. This often results in underfitting as the models strive to minimize overall prediction errors. Additionally (ii), the spatial distribution of snow depth is inherently scale-dependent [60, 66], meaning a sensor’s spatial resolution is critical in determining captured variability. Our training samples, derived from the ICESat-2 ATL08 elevation product, correspond to an area of ca. 20 m × 15 m, which is better than previous study cases but still averages out for finer-scale variability. ICESat-2’s sparse spatial sampling pattern with several kilometres gap between sampled snow depth profiles, combined with a 92-day revisit period of the satellite and possible cloud cover leads to substantial gaps in seasonal data coverage. Data points with high snow depth measurements are thus likely to be relatively few. ICESat-2’s elevation data, derived from a ground-finding algorithm that use a probability distribution function (PDF) of reflected photons [73], can vary in quality depending on the terrain. In rugged terrain where snow depth is typically the deepest, noise or unbalanced sampling can lead to misrepresentation of extreme values.
@@
The shortcomings of the model and data sample could be mitigated if model predictors were able to fully capture spatial variability of snow depth and its underlying physical processes (iii), which is not the case. ERA5 Land input snow depth values may provide incorrect prior information as the dataset seems to contain interpolation effects that are possibly a downscaling artefact from the production of ERA5 Land from the coarser ERA5 data [35]. For example, snow depth values are much higher for ERA5 Land grid cells on or adjacent to large glaciers like the Jostedalsbreen ice cap in Southern Norway, and lower for grid cells on or adjacent to large water area. Consequently, although we have excluded all glaciers and lakes, neighbouring ERA5 Land grid cells still contain these bias effects. There, the predicted subgrid variability is not solely a reflection of inherent snow conditions but is significantly influenced by this spatially-dependent bias that the model needs to detect and correct for based on spatial information, i.e. coordinate features (N, E). This highlights the importance of dense, localized sampling of ICESat-2 snow depths, which may be challenging for smaller study areas and for lower latitudes, given ICESat-2’s sparse sampling pattern. By expanding the training area, as done in this study, the model benefits from more samples but might also learn a more generalized, averaged representation of snow depth.
Among the predictor features, elevation emerges as the most significant, as expected from the temperature–elevation relationship, orographic and precipitation shadowing effects [66, 80] where also aspect plays an important role. The TPI at a 90-meter scale is a strong predictor (Figure 9) and corresponds to the size of terrain features in our study area. However, these predictors represent general patterns only. They are not sufficient to represent the locally snow depth variability. Here, the cumulative effect of multiple snow transport events over an entire winter season plays a crucial role. By refining the wind-aspect factor [3, 17] into new cumulative values to describe wind redistribution (i.e. accumulation or erosion), our model is able to explain a considerable amount of the observed preferential snow deposition (Figure 9 g). Erosion is not well captured (Figure 9 f) by the downscaling model. There is potential to improve wind-related predictors, as the spatial-temporal resolution of wind fields from ERA5 Land monthly data is likely insufficient for capturing complex wind-topography interactions. Downscaling of wind fields [21, 93] could enhance the model’s ability to capture wind-driven variability. Additionally, the model does not account for the impact of wind on the energy balance at the snow-atmosphere interface, involving sensible and latent heat exchanges [65], a factor crucial to the dynamics of a snowpack. The choice of predictors in this study is based on our expertise and corresponds to commonly used predictors in statistical modelling of snow distribution. This choice inherently introduces bias as the predictors are not able to explain the entire variability. Future research could benefit from incorporating more features that are related to physical snow processes or the integration of snow melt-out information from fractional snow cover data from optical satellite imagery [24] commonly used in snow data assimilation schemes with process-based models [56, 57]. Used as an exploratory tool, the model might assist in learning what processes are at play and how to best represent them statistically, ideally leading to a simplification of the workflow.
@@
The calibration step uses quantile mapping based on the observed cumulative snow depth distribution function during 2009 peak snow conditions in the Hardangervidda area. Our selected control points cover the entire study area (~1.5 points per square kilometre), ensuring a comprehensive correction across all regions analyzed. While greatly improving our results, especially for the reproduction of extremes values, the need of a scaling calibration is a disadvantage for an extrapolation of our method to global applications. The slight underestimation of snow depths for the presented western sub-region in the validation area (Figure 6) suggests that the single scaling function obtained from the Hardangervidda area does not fit the sub-region perfectly, and a locally adjusted scaling might result in a better fit.
Future research is needed to determine optimal training/model area sizes or transferable scaling calibration functions for areas with different topography and climate. However, extensive, other publicly available snow depth reference datasets like the Hardangervidda ALS dataset are currently not available for Norway and are generally very rare. In the western United States, the Airborne Snow Observatory data [ASO; 78] includes multi-temporal snow depth maps of several large catchments that could be suitable to examine the local dependency of our downscaling approach and further develop the method. Similarly, the lidar dataset collected in Northern Alaska by [80] could be used to train for an Arctic snowpack. As an alternative to using extensive validation data for model result calibration, future research could explore the use of small representative snow depth distributions such as strategically chosen, smaller snow depth maps from drone-based surveys, data from snow courses, or snow depth measurements from meteorological stations.
@@ 5.3 Validation
Snow depth distribution in alpine environment is influenced by distinct processes occurring at various spatial and temporal scales. Our model demonstrates high fidelity of snow distribution patterns even at the microscale (Figure 6 c), though with lower performance metrics compared to the mesoscale. Our findings complement the insights from Mott et al. [66], which recommend a resolution finer than 50 meters to capture the spatial variability inherent to wind-driven processes. For example, Trujillo et al. [95], Mott et al. [64], Mendoza et al. [61] report a distinct ’scale break’ at around 100 m, with a stronger spatial autocorrelation of snow depths below the scale break than beyond. Such spatial correlation can be attributed to the wavelength of wind-driven snow accumulation features [66]. The ALS data and our model output identified the presence of scale breaks and dependence on terrain features (95 m accounting for 79% variance for ALS data and 108 m for 60% variance for our model, Figure 10), demonstrating the effectiveness of our scheme in capturing a high level of spatial details and the impact of landscape patterns. Notably, we also found a scale break at 1,300 m by fitting the empirical variograms at two different ranges, which corresponds to the wavelength of ridges in this area (Figure 10). These scale breaks underscore the critical resolution and minimal survey scope required to accurately capture spatial variability over the hillslope. Such insights are pivotal for designing snow surveys or evaluating snow models/products.
RMSE or $R2$ are standard metrics for evaluating regression model performance, focusing primarily on overall accuracy and fit. However, they may not adequately capture the variability and structural patterns in snow depth distribution, especially at the microscale. Figure 6 and 7 show high signal fidelity (high Spearman rank correlation) however with poor $R2$. To enhance the evaluation of regression outcomes for snow depth subgrid variability, it is essential to consider the scales involved and incorporate additional metrics. As an alternative to the Spearman rank correlation used in this study, structural similarity index measure (SSIM, Parr et al. 80) might be a suitable metric to capture relative similarities.
@@ 5.4 Application
Current limitations of ICESat-2 data, including the month-long data availability lag and significant spatial and temporal gaps, restrict its utility [15]. Spatio-temporal propagation of ICESat-2 snow depths by use of modelling is thus the logical next step to produce timeseries of snow depth maps and make this data useful for wider applications. Hereby, standard statistical models that solely depend on topographic features for snow spatial distribution fall short on temporal variability. Our workflow resolves many of these constraints by incorporating temporal variability from ERA5 Land. Other emerging approaches to create snow depth maps from ICESat-2 data include combining ICESat-2 snow depths with snow cover information and a process-based model within a data assimilation framework (Mazzolini et al. 57, using higher-resolution snow depths from ATL03 data) or a combination of deep learning and probabilistic data assimilation methods (Guidicelli et al. 29, using snow depth tracks synthesised from snow depth maps to mimic ICESat-2 snow depths). Both approaches yield promising results and also use ERA5 Land data as forcing, thus in principle globally applicable, but focusing on smaller catchments and coarser spatial resolution than presented in this work. Future research could focus on comparing or merging these different approaches.
Notably, we find that the performance of aggregated downscaled snow depth is not sensitive to the absolute precision of the reference DEMs (Supplementary Section D), indicating that our approach is not limited to areas with high-resolution DEMs but can be applied using global DEMs in treeless conditions globally. However, careful consideration is required when applying this workflow to new study areas or applications:
  1. 1.
    Global DEMs like the Copernicus GLO-30 or FABDEM are a patchwork of data from satellite-derived DEMs and local/national data sources and may not have equally good quality in different regions,
  2. 2.
    Care has to be taken to ensure including enough ICESat-2 samples in/around the target area. Areas further south will have larger coverage gaps due to ICESat-2’s sampling pattern that is densest at the poles. Upcoming ICESat-2 data expected for several more years will mitigate this
  3. 3.
    Applying calibration with the most representative field survey
  4. 4.
    Accounting for regional climatic variability that some area are heavily influenced by specific weather events or patterns.
@@
The workflow described herein is applicable to a range of applications in need to assess snow distribution patterns where sparse snow observations exist. For example, our workflow can (1) mitigate data gaps in remote areas spatially and temporally, e.g. by interpolating and extrapolating weather station observations (Figure 6, 7), (2) help to plan and reduce cost of new snow surveys (e.g coordinating surveys with ICESat-2 coverage), (3) improve local estimates of snow water equivalent in large alpine watershed, and (4) be combined with a classical snow modeling approach. Suppose a snow model provides not just average value snow depth, but also a probability density distribution representing the subgrid variability [26], such a curve can then be translated into a snow depth map using our snow downscaling scheme.
In the current implementation, we observe some limitations in extrapolating the method in time, or simulating timeseries for a single point. First, our current implementation assumes stationarity of ERA5 Land data. For instance, [43] found discontinuities in the ERA5 Land snow products around 2004 due to the introduction of new satellite products in the assimilation scheme. Other stationarity issues may arise from the method used to generate ERA5 Land with successive spin-up periods [1949, 1981, 2001; 86] or due to changed climate conditions in the future. An alternative to expand the temporal extrapolation and improve stationarity issues would be to employ an energy-mass balance snow model like Crocus or Snowpack [99, 2] using ERA5 as forcing. Such snow reanalysis approaches were found to be better at estimating snow water equivalent (SWE) globally than methods based on passive microwave observations [63]. Our statistical approach performs well compared to the seNorge product, but there is currently no physically-based snow reanalysis available at a comparable spatial resolution in Norway. Any subsequent application of the proposed approach will require careful assessment of uncertainties and the quality of input data, including ERA5 Land stationarity. Second, simulating snow depth at the point scale as shown in section 4.3 shows poor agreement with station measurements (Table 2), and the application of a simple correction factor to ERA5 Land timeseries would likely yield better results. This is partly due to various local processes that may affect snow distribution not captured by the model (e.g. localised wind field in respect to terrain, radiation, etc.) or locally disturbing the weather station, but also the model capabilities and design with mostly spatially constant input features. Further studies may consider adding predictor features that consider time explicitly or in a summarised way over the course of the snow season, such as a radiation or heat index [12] or annual snow melt-out dates from fractional snow cover data.
@@
Given careful calibration and information on snow density, the presented cost-efficient subgrid parameterization for snow depth could be used to estimate SWE or to correct precipitation/snowfall bias in snow models [25], to eventually serving as input for discharge modeling [34]. Microscale snow depth maps could benefit studies of ecosystems in snow-covered regions, from habitat availability for wildlife [51] to plant phenology under the snowpack [75]. Snow depth parametrizations/distributions at the mesoscale for large areas or detailed snow depth maps at the microscale are crucial for local permafrost studies [26].
@@ 6 Conclusion
This study introduces a workflow for snow depth retrieval from ICESat-2 ATL08 and DEM data to downscale ERA5 Land snow depth data using XGBoost tree-structure machine learning models. The two datasets, ICESat-2 and ERA5 Land, have complementary resolutions in space and time that allow for the generation of accurate monthly snow depth maps at the hillslope scale. Hereby, ERA5 Land data primarily provides the temporal variability, and ICESat-2 the spatial variability in snow depths, propagated in space using terrain features and other relevant predictors to train the downscaling model. Advanced bias correction and calibration are part of the workflow to address inherent systematic errors present in the data and correct for residual bias.
@@
    There are few snow depth observations available in remote areas, and to this day, no inexpensive ways to map small-scale variability exist. Here, ICESat-2 ATL08 data as presented in our workflow of co-registration and bias-correction stands out as a valuable data source.
@@
    The downscaling-calibration scheme’s performance to predict peak snow for 6×2 ALS flight strips in the Hardangervidda is very good at mesoscale (100 x 500 m, $R2$ values ranging from 0.74 to 0.88). At microscale (10 m), the spatial snow depth pattern is captured very well but absolute values are less so. The model is also able to represent the spatial pattern of snow melt-out during late spring as visible from snow cover satellite data.
@@
    Post-calibration of downscaling model results is currently necessary to compensate for the model’s under-representation of extreme values. This under-representation is likely caused by inherent model behaviour (tending towards average values), remaining bias and the nature of the training data. Biasing factors include persistent snow patches on supposedly snow-free reference DEMs, simplified representation of wind processes, under-sampling and underestimation of high snow depths in ICESat-2 data, and temporal non-stationarity in ERA5 Land data. Future research and more adaptable calibration methods for varied scenarios may improve results and remove the need of post-calibration.
@@
This work was supported by the SNOWDEPTH project, Research Council of Norway (Grant number 325519) and the research project ClimLAND, an EEA/EU collaboration grant between Norway and Romania 2014-2021, project code RO-NO-2019-0415,1290 contract no. 30/2020. It is a contribution to the strategic research initiative LATICE (University of Oslo, project UiO/GEO103920) and also got some support from the French Ministry of Ecological Transition (convention 261168).